OpenStudy (anonymous):

((8x)/(3x-4))((x^2-4)/(2x+4)) Where do I even begin tryin to multiply these?

OpenStudy (anonymous):

do i multiply across then simplify or simplify first? or is any of what Im sayin right at all?

OpenStudy (anonymous):

\[\frac{8x}{3x-4}\times \frac{x^2-4}{2x+4}\] step one is to factor as \[\frac{8x(x-2)(x+2)}{(3x-4)2(x+2)}\] step two is to cancel the common factor of x+2 to get \[\frac{8x(x-2)}{2(3x-4)}\] and i think that is the best you can do unless you want to multiply out in the numerator and denominator